OpenStudy (anonymous):

Questions on work (I am lost in this so any help would be appreciated): 1. A frictionless 1kg cart is initially at rest. If you push the cart 1m on level ground with a constant horizontal force of 1N, what is its final speed? 2. What is the work needed to lift a 1kg box 1m in the air? 3. If you use a two pulley system to lift a 1kg box, so that the tension need to lift the box is only half of the box's weight, what is the work needed to lift a 1kg box 1m in the air? 4. What is the power needed to drive a 1000kg car up a 30° incline at a constant speed of 10m/s? Ignore air resistance

OpenStudy (anonymous):

5. An 8nm thick cell membrane maintains an electric potential of 100mV between the cytoplasm (top) and cell exterior (bottom). This acts like a battery to power ATP synthase. If it takes 1/10th of an electron-Volt (0.1eV = 1.6x10-20J, energy) for the electric gradient to pull a proton through the membrane, what is the average electric force on a proton as it is pulled through the membrane? (Is ΔElectric PE + or – ?) A) 0 N B) 2x10-12 N C) 5x1010 N D) not enough information

OpenStudy (anonymous):

at 1, u have to mention the time

OpenStudy (anonymous):

@Arnab09 it wasn't given

OpenStudy (anonymous):

Actually I think 1 is sqrt2 m/s and 2 is 9.8 J so i could cross those off my list

OpenStudy (anonymous):

hey u der i can help u for 1 qn we have u=0 v=? a=1 f=1 m=1 s=1 now we can fint v by newtons third law v^2=u^2 + 2as

OpenStudy (anonymous):

v^2=2 v=sqrt2

OpenStudy (anonymous):

v^2=2 v=sqrt2 v=1.414 m/sec

OpenStudy (anonymous):

F=mg and theta will be 90 degree u knw work =FScostheta work=10x1 x0 work=0

OpenStudy (anonymous):

 this is the figure of 3 qn